OpenStudy (barrera1):

The distance you travel varies directly with the amount of gas you use. You can travel 300 miles on 12 gallons of gas.

OpenStudy (barrera1):

Question: a. You want to know how much gas it will take to for your entire road trip that is 750 miles. Write a proportion to solve this equation. b. How many gallons will it take? c. If gas is $3.75 a gallon, how much will it cost in gas to go on your road trip? d. What is the cost per mile?

OpenStudy (mathmale):

Hint: Start with the ratio\[\frac{ 300~miles }{ 12~gallons }\]

OpenStudy (barrera1):

I divide that right? @mathmale

OpenStudy (anonymous):

Leave it in that form and set up an equivalent ratio

OpenStudy (barrera1):

I don't understand, What do you mean?

OpenStudy (anonymous):

You know youre going to travel 750 miles but you dont know how much gas you need right

OpenStudy (barrera1):

oh okay okay

OpenStudy (anonymous):

You know you can travel 300 miles on 12 gallonsof gas which when put in a ratio is 300/12. How would you put 750 and x in a ratio

OpenStudy (anonymous):

750 miles/X gallons of gas

OpenStudy (anonymous):

Now you need a proportion.

OpenStudy (barrera1):

okay

OpenStudy (anonymous):

300 miles/12 gallons of gas =750 miles/ X gallons of gas

OpenStudy (anonymous):

Do you know how to cross multiply?

OpenStudy (barrera1):

yes

OpenStudy (anonymous):

Why don't you go ahead and do that

OpenStudy (barrera1):

I'm confused at this point what am I cross multiplying?

OpenStudy (anonymous):

You have the fraction 300/12 = 750/X

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

750*12 = 300*X

OpenStudy (barrera1):

30?

OpenStudy (anonymous):

Yes, very good

OpenStudy (anonymous):

What does that 30 represent?

OpenStudy (barrera1):

gallons?

OpenStudy (anonymous):

Yeah, you need 30 gallons of gas to make a 750 mile road trip. Now it wants you to calculate the price of 30 gallons of gas.

OpenStudy (anonymous):

Can you do the rest by yourself?

OpenStudy (barrera1):

um yeah just can you help me on D? that's all I need help on then I'll give you a medal is that fine?

OpenStudy (anonymous):

ok give me a sec, my browser is acting up

OpenStudy (anonymous):

So its asking you how much it costs for each mile. How do you think you could figure that out?

OpenStudy (anonymous):

Your total cost is 30 * 3.75 right?

OpenStudy (barrera1):

112.5?

OpenStudy (anonymous):

Yes thats how much the entire trip cost. 750 miles cost you 112.50. How much would one mile cost?

OpenStudy (barrera1):

I divide now?

OpenStudy (anonymous):

Think about it, if you divided the total cost by the total number of miles what would that give you

OpenStudy (anonymous):

It would give you the cost for one mile

OpenStudy (anonymous):

Did you divide them?

OpenStudy (anonymous):

112.50 Total cost/750 miles = cost/1 mile

OpenStudy (anonymous):

What did you get Barrera?

OpenStudy (barrera1):

a lot of 6's lol

OpenStudy (barrera1):

so no dividing and not multiplying right?

OpenStudy (anonymous):

You divided 112.50 by 750?

OpenStudy (barrera1):

0.15? I divided wrong oops

OpenStudy (anonymous):

Yeah, it cost 15 cents a mile

OpenStudy (barrera1):

oh okay thanks for the help

OpenStudy (anonymous):

no problem